if cos x= 3cos y then 2tan (y-x)/2 is equal to 
in form of cot
 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
 
write it as cosx/cosy = 3/1
now apply componendo and dividendo
(cosx+cosy)/(cosx – cosy)= (3+1)/(3-1)
now apply formulae for cosx+cosy and cosx – cosy
2tan(y – x)/2= cot(x+y)/2
 
Hope it helps